# BranchGRPO: Stable and Efficient GRPO with Structured Branching in Diffusion Models

### v4

@@ 4.5 Scaling with Branch
DanceGRPO scales poorly: generating 81 rollout samples requires over 2400s per iteration, whereas BranchGRPO achieves the same scale in only 680s, making large-scale scaling feasible. We investigate two settings: scaling the branch factor (K=2,3,4K{=}2,3,4 yielding 16, 81, and 256 leaves) and scaling the number of branching steps ($3,4,5$ splits yielding 8, 16, and 32 leaves under K=2K{=}2).
@@
Figure 6: Impact of scaling branch rollouts in BranchGRPO. Larger branch factors (a) and more branching steps (b) consistently improve reward, while iteration time (c) increases with leaves.
Figure 7: Video generation results on WanX-1.3B. Left: qualitative frame comparisons across three settings. Right: reward curves showing faster convergence and higher final rewards with BranchGRPO compared to DanceGRPO.
@@ 4.6 Video Generation Results
We further evaluate BranchGRPO on video generation using WanX-1.3B (Wan et al., 2025), with the Video-Align motion quality (Liu et al., 2025b) as reward.
@@ 5 Conclusion
We presented BranchGRPO, a tree-structured reinforcement learning framework that overcomes inefficiency and sparse credit assignment in GRPO by introducing branching rollouts with shared prefixes, reward fusion with depth-wise advantages, and lightweight pruning. Experiments on image and video generation show that BranchGRPO delivers faster, more stable, and higher-quality alignment, establishing structured branching as a practical and scalable path for RLHF in diffusion and flow models.
Appendix A Appendix
@@
  • Section B: Hyperparameter settings used in all experiments.
  • Section C: Theoretical analysis, including proofs for branch noise construction, reward fusion, and variance reduction.
  • Section E: Additional experiments, including ablations, more text-to-image and image-to-video results.
  • Failure Cases: Qualitative examples where BranchGRPO fails, highlighting current limitations.
Appendix B Hyperparameter Settings
Table 2 summarizes the detailed hyperparameter configuration used in our experiments. All hyperparameters are kept identical across all methods, including DanceGRPO and MixGRPO, to ensure a fair comparison.
@@ Appendix B Hyperparameter Settings
Appendix C Theoretical Analysis
C.1 Branch Noise Construction and Boundary Distribution Preservation
@@
C.2 Reward Fusion: Unbiasedness and Variance Reduction
@@
C.3 Depth-wise Baseline (Control Variates)
@@
C.4 Continuous Rewards and Concentration
@@
Appendix D Experiment Setup Details
This section provides extended details of our experimental setup, complementing the summary in Section 4.3 of the main paper. We describe dataset splits, model backbones, branching configurations, pruning strategies, and evaluation protocols.
Datasets.
We conduct experiments on HPDv2.1 (Wu et al., 2023), which contains 103k training prompts and 400 balanced test prompts across diverse styles. For video generation, we adopt WanX-1.3B (Wan et al., 2025) on the image-to-video (I2V) task, following the standard evaluation with Video-Align (Liu et al., 2025b).
Model Backbone.
We use FLUX.1-Dev (batifol2025flux) as the text-to-image backbone. All experiments are initialized from the same pretrained checkpoint. For video tasks, we build on the WanX-1.3B model.
Branching Configurations.
We set default tree depth to $d=20$ and branching factor $K=2$, yielding 16 leaves per rollout before pruning. Branching steps $\mathcal{B}$ follow three schedules: Dense $(0,3,6,9)$, Mixed $(0,4,8,12)$, and Sparse $(0,5,10,15)$. We sweep correlation $s\in\{0,1,2,4,8\}$ to study its effect. All SDE steps are retained in forward rollouts; pruning is only applied after the final branching depth, while width pruning selects among leaves at the last split. For fair comparison, depth and width pruning drop the same number of gradient steps.
Evaluation Metrics.
We report efficiency metrics including per-iteration wall-clock time, effective number of function evaluations (NFE), and GPU-hours. For quality, we use HPS-v2.1 (Wu et al., 2023), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2023). For video generation, we use Video-Align (Liu et al., 2025b). Additional metrics (KID (Bińkowski et al., 2018), MMD2 (Gretton et al., 2012), IntraFID (kynkaanniemi2019improved), CLIP-based metrics (radford2021clip)) are reported in the diversity analysis.
Appendix E Additional Experiments
E.1 More quantitative result
@@
Uniform Fusion 13.68 13.68 493 0.361 0.229 1.218 0.368
Path-Weighted Fusion 13.68 13.68 493 0.363 0.229 1.233 0.374
@@
Width Pruning 13.68 8.625 314 0.364 0.230 1.300 0.374
Depth Pruning 13.68 8.625 314 0.369 0.231 1.319 0.381
E.2 More Text2Image Results
@@
E.3 More Image2Video Results
@@ E.4 Failure Cases
Appendix F Discussion, Limitations, and Future Work
@@ Discussion.
Limitations.
First, our study is limited to image alignment on HPDv2.1 and video alignment on WanX-1.3B. Although these benchmarks are representative, broader validation on more diverse datasets and larger-scale video models is necessary. Second, our current design relies on pre-specified branching schedules and fixed correlation strengths; while effective, this may underexplore adaptive structures that respond to sample difficulty or reward uncertainty. Finally, our pruning strategies reduce gradient cost but still require full forward rollouts, leaving memory and bandwidth overhead partially unaddressed.

### v5

@@ 4.5 Scaling with Branch
DanceGRPO scales poorly: one GRPO training step with 81 rollout samples takes over 3500s, whereas BranchGRPO achieves the same scale in only 680s, making large-scale scaling feasible. We investigate two settings: scaling the branch factor (K=2,3,4K{=}2,3,4 yielding 16, 81, and 256 leaves) and scaling the number of branching steps ($3,4,5$ splits yielding 8, 16, and 32 leaves under K=2K{=}2).
@@
Figure 6: Impact of scaling branch rollouts in BranchGRPO. Larger branch factors (a) and more branching steps (b) consistently improve reward.(c) Combining both dimensions, larger group sizes yield consistent reward gains, following a clear scaling law.
Figure 7: Video generation results on Wan2.1-1.3B. Left: qualitative frame comparisons across three settings. Right: reward curves showing faster convergence and higher final rewards with BranchGRPO compared to DanceGRPO.
@@ 4.6 Video Generation Results
We further evaluate BranchGRPO on video generation using Wan2.1-1.3B (Wan et al., 2025), with the Video-Align’s motion quality (Liu et al., 2025b) as reward.
@@ 5 Conclusion
We introduced BranchGRPO, a tree-structured GRPO paradigm that replaces sequential rollouts with prefix-sharing branching and depth-wise reward fusion, augmented by lightweight pruning for compute reallocation. Across image and video generation, BranchGRPO yields faster convergence, more stable training, and higher final alignment quality under matched budgets. These results establish structured branching as a practical, scalable path for RLHF in diffusion and flow models, and we further verify a scaling trend in which larger group sizes consistently improve performance.
Ethics Statement
This work focuses on improving optimization efficiency and stability for diffusion/flow-based generative models. Our experiments rely on publicly available prompt sets and reward models cited in the paper. No personally identifiable information, human subjects data, or sensitive attributes were collected or annotated by the authors. We followed the licenses and usage terms of all datasets and third-party models; when applicable, we restricted use to non-commercial research. Given that generative models may amplify biases in reward models or datasets, we report failure cases and ablations that probe stability and variance, and we avoid deploying the models in safety-critical settings. We will release code and configuration files to facilitate scrutiny and responsible reuse.
Reproducibility Statement
We release anonymized training and evaluation code, configuration files, and exact hyperparameters in the supplementary materials. All experiments can be reproduced using the provided scripts, seeds, and environment specifications.
Appendix
@@
  • Section A: Hyperparameter settings used in all experiments.
  • Section B: Theoretical analysis, including proofs for branch noise construction, reward fusion, and variance reduction.
  • Section C: Additional experiments, including ablations, more text-to-image and image-to-video results. Failure Cases: Qualitative examples where BranchGRPO fails, highlighting current limitations.
  • Section D: Discussion and future work.
LLM Usage
An LLM was used only for language polishing of the manuscript (e.g., grammar, wording, and minor clarity edits). No experimental results were generated, modified, or selected by the LLM; all technical content, claims, ideas, and analyses are the authors’ own. No confidential or non-public data were shared with the LLM. The authors take full responsibility for the final content.
Appendix A Hyperparameter Settings
Table 2 summarizes the detailed hyperparameter configuration used in our experiments. All hyperparameters are kept identical across all methods, including DanceGRPO and MixGRPO, to ensure a fair comparison. For depth pruning and hybrid-ODE-SDE, we follow the design of MixGRPO and adopt a sliding window of size 4, which shifts one step deeper every 30 iterations.
@@ Appendix A Hyperparameter Settings
Appendix B Theoretical Analysis
B.1 Branch Noise Construction and Boundary Distribution Preservation
@@
B.2 Reward Fusion: Unbiasedness and Variance Reduction
@@
B.3 Depth-wise Baseline (Control Variates)
@@
B.4 Continuous Rewards and Concentration
@@
Appendix C Additional Experiments
C.1 More quantitative result
@@
Uniform Fusion 13.68 13.68 493 0.361 0.229 1.595 0.368
Path-Weighted Fusion 13.68 13.68 493 0.363 0.229 1.603 0.374
@@
Width Pruning 13.68 8.625 314 0.364 0.231 1.609 0.374
Depth Pruning 13.68 8.625 314 0.369 0.235 1.625 0.381
C.2 More Text2Image Results
@@
C.3 More Image2Video Results
Table 4: Video evaluation on vBench. We report mean values on 500 samples.
Method Aesthetic Background Dynamic Imaging Motion Iteration
Quality Consistency Degree Quality Smoothness Time (s)
Base Model 0.5206 0.9588 0.5150 71.92 0.9784 -
DanceGRPO 0.5178 0.9647 0.4992 71.94 0.9899 1352
BranchGRPO 0.5190 0.9659 0.5000 71.94 0.9912 493
@@ C.4 Failure Cases
Appendix D Discussion and Future Work